# Seeing Beyond Haze: Generative Nighttime Image Dehazing

### v1

@@ B.2 Controllable Generative Dehazing
Figure S2: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S3: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S4: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S5: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S6: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S7: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure S8: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.

### v2

@@ Appendix F Societal Impact
Figure 8: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 9: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 10: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 11: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 12: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 13: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
@@
Figure 14: Qualitative results from DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
Figure 15: Ablation study on real-world haze datasets. “Pre-trained” denotes the output of our pre-trained dehazing model. “Ours w/o supp” uses only our generative dehazing priors. “Ours” combines both our generative priors and supplementary components.